OpenStudy (diamondanon2):

Find the area of the triangle ABC given angle B=45 degrees, side a = 10, b = sqrt2

jimthompson5910 (jim_thompson5910):

step 1) use the law of sines to find angle A sin(A)/a = sin(B)/b sin(A)/10 = sin(45)/sqrt(2) solve for A step 2) use angle A and B to find angle C A+B+C = 180 degrees C = 180-A-B step 3) use sides 'a' and 'b', along with angle C, in the area formula area = (1/2)*a*b*sin(C)

OpenStudy (diamondanon2):

Answer choices are 5 5sqrt2 10 10sqrt2

OpenStudy (diamondanon2):

Thanks a bunch! :)

jimthompson5910 (jim_thompson5910):

hmm there's probably a typo in the problem when I tried to solve sin(A)/10 = sin(45)/sqrt(2) I ended up with sin(A) = 5 but that has no solutions

OpenStudy (diamondanon2):

Yeah that's what I got too

OpenStudy (diamondanon2):

Is there a way to find the third side to calculate the area?

jimthompson5910 (jim_thompson5910):

law of cosines c^2 = a^2 + b^2 - 2*a*b*cos(C) but you'd need to know the angle C

OpenStudy (diamondanon2):

Hmm them so this is unsolvable the way it is?

OpenStudy (diamondanon2):

*then

jimthompson5910 (jim_thompson5910):

I just used geogebra to try to construct this triangle visually. I'm seeing that it's not possible with these sides and angle can I see a screenshot of the full problem?

OpenStudy (diamondanon2):

Yeah I'm using a mobile device though could I send it to you another way?

jimthompson5910 (jim_thompson5910):

sure, send it however you want

OpenStudy (diamondanon2):

Sent it

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

we know 2 sides and the angle between them so we use the SAS triangle area formula area = (1/2)*(side1)*(side2)*sin(angle between sides) area = (1/2)*(a)*(c)*sin(B) area = (1/2)*(10)*(sqrt(2))*sin(45) I'll let you finish up

jimthompson5910 (jim_thompson5910):

let me know if you have any questions @DiamondAnon2